# Membership Inference Attack with Partial Features

### v1

@@ V-E Ablation Study
V-F Impact of Feature Importance on Attack Performance
Since models perform inference based on feature values and different features contribute unequally to the model output, we investigate how the importance of known features influences attack performance. In this section, we use SHAP [42] to quantify the contribution of each feature to the prediction of the target model. SHAP is a widely used method for interpreting matchine learning models by assigning an importance value to each feature. It is based on the concept of Shapley value from game theory. Features that contribute more to the output are assigned higher importance values. We evaluate the attack performance under different known feature proportions, using two selection strategies:
@@ V-F Impact of Feature Importance on Attack Performance
  • When the known features are those with high SHAP importance, the attack achieves significantly better performance compared to when unimportant features are known. This is likely because important features drive the model prediction, while unimportant known features lead to vague optimization directions and weak gradient signals.
  • When the percentage of known features is small enough, using important features as known features outperforms random selection. We attribute this to the fact that important features are more likely to activate memorized patterns in the model, which helps ensure the effectiveness of the reconstruction even when only a small number of such features are available.
  • When the percentage of known features exceeds a certain threshold, using only important features as known ones may yield lower performance compared to random selection. This is because when all important features are already known and the remaining few features are non‑important, the unknown features have little influence on the loss. As a result, feature reconstruction can hardly yield additional useful membership information.
@@
Figure 9: Impact of known feature importance on attack performance (AUC) across different datasets. Solid lines represent scenarios where known features are the most important ones (K=IF), while dashed lines indicate scenarios where the least important features are known (K=UF). We denote the case where only one feature is unknown as “loo”.
@@ V-G Case Study
To simulate this scenario, we use the UCI-Diabetes dataset. We select 10 features for model training and designate race, age, and gender, common demographic attributes easily accessible to an attacker, as known features. Other experimental settings remain consistent with previous sections. The attack results are shown in Table V, demonstrating that our method can effectively compromise individual privacy in practical settings.
TABLE V: AUC performance of case study.
@@ VI Discussion
VI-A Defenses
As with other inference attacks, overfitting plays a key role in the success of the attack. Therefore, any technique that mitigates overfitting, such as early stopping [43], dropout [44], and other regularization methods, can serve as a potential defense.
Differential privacy[45], as a provable privacy-preserving framework, also serves as a viable defense strategy. By introducing randomness into the system, it effectively reduces the risk of inferring the presence of any individual in the whole data group. Abadi et al.[46] introduced differential privacy into the field of machine learning by proposing the DP-SGD training framework, which provides formal privacy guarantees during model optimization. This approach has been widely adopted as a defense against membership inference attack.
VI-B Relationship to Other Attacks
Relationship to Property Existence Attack: Property existence attack[18], as a special case of distribution inference attack[18, 47], aims to determine whether a given feature combination exists in the training set. The only difference from our attack lies in whether the attacker has access to the unknown features. Clearly, an attacker capable of performing PFMI can easily carry out a property existence attack, but it is difficult for a property existence attacker to perform PFMI because of the absence of some features.
Relationship to Model Inversion Attack: A common premise of existing model inversion attacks is that the adversary knows the target to be present in the training set[30, 31, 32, 33, 34, 35]. Even though the membership status of the target is unknown in the attack scenario considered in this paper, the first stage of our attack can still be regarded as a form of model inversion attack. Leveraging existing model inversion techniques thus represents a feasible approach for improving performance.
VI-C Limitation and Future Work
While our proposed method has demonstrated promising attack performance, several limitations remain.
First, the attack presented in this paper is based on a white-box setting, which may limit its applicability in certain scenarios. To bridge this gap, one possible direction is to explore gradient approximation techniques[48, 49], which allow attackers to approximate gradients without full model access. Additionally, model stealing methods[50, 51] could be leveraged to construct a white-box surrogate of the target model, thereby enabling the application of our attack in a black-box context.
Secondly, when the percentage of known features becomes too large, the attack performance tends to degrade. Based on our experimental results, one potential solution is to selectively discard some known features during inference—particularly those that, while somewhat influential to the prediction of the model, are not essential for preserving the representativeness of the sample with respect to the underlying individual. This strategy may help mitigate performance drops and enhance the overall effectiveness of the attack.
Furthermore, due to the missing feature values, the mapping between the known features and the label may be one-to-many. This paper assumes that the attacker can obtain at least one valid label, which is a realistic assumption in practice. However, exploring label-free attacks could further enhance the capability of the attacker in more extreme scenarios and is therefore a worthwhile direction for future research.
@@ VII Conclusion
In this paper, we study membership inference attack under the setting where some feature values are missing. To address this scenario, we introduce a practical two-stage attack framework named MRAD. Our framework is compatible with arbitrary anomaly detection methods, enabling attackers to flexibly conduct attacks under various knowledge settings.
We conduct extensive experiments on multiple datasets to evaluate the effectiveness of our method. The results show that our framework consistently delivers strong performance across a wide range of known feature percentages. In most datasets, once the known feature ratio exceeds 50%, the attack begins to produce reliable results. Remarkably, even when 40% of the features are missing, the attack still achieves an AUC of 0.6. In addition, we use SHAP to measure feature importance and experimentally confirm that the importance of the known features has a significant impact on the attack’s performance. Finally, through a case study, we simulate a real‑world application scenario to further highlight the practical relevance of our attack method.
Our findings demonstrate that membership inference attacks can be mounted using publicly available partial data, even when a sample’s private features are protected. This highlights a new dimension in privacy risk assessment.
Acknowledgment
The authors would like to thank…

### v2

@@ V-D Ablation Study
V-E Impact of Feature Importance on Attack Performance
Since models perform inference based on feature values and different features contribute unequally to the model output, we investigate how the importance of known features influences attack performance. In this section, we use SHAP [NIPS2017_7062] to quantify the contribution of each feature to the prediction of the target model. SHAP is a widely used method for interpreting matchine learning models by assigning an importance value to each feature. It is based on the concept of Shapley value from game theory. Features that contribute more to the output are assigned higher importance values. We evaluate the attack performance under different known feature proportions, using two selection strategies:
@@ V-E Impact of Feature Importance on Attack Performance
  • When the known features align with those that have high SHAP importance, the attack performs much better than when only low-importance features are available. This is because important features have a strong influence on the model’s predictions, whereas unimportant features contribute very little. As a result, those unimportant features provide only weak gradient signals and offer less meaningful guidance for the optimization process of the first stage.
  • By comparing Figure 3 and Figure 11, we observe that when a large portion of the features is known, using only high-importance features as the known ones can actually lead to worse performance than selecting the known features at random, particularly for STL-10 and Epsilon. This happens because once most important features are already known, the remaining unknown features are largely non-essential and contribute very little to the loss. In other words, the values of these remaining features can vary almost arbitrarily without meaningfully affecting the model’s prediction, leaving the attack with too little membership-relevant signal to exploit.
These results offer insights for both launching and defending against such attacks. For attackers, achieving strong performance requires keeping the number of important known features within a reasonable range—having too few or too many can both lead to degraded attack effectiveness. For defenders, instead of protecting a large number of features, it may be more effective to prioritize the protection of high-importance features, as these features contribute most to the memorization behavior of the model.
@@
Figure 11: Impact of known feature importance on attack performance (AUC) across different datasets. Dashed lines represent scenarios where known features are the most important ones (K=IF), while solid lines indicate scenarios where the least important features are known (K=UF). We denote the case where only one feature is unknown as “loo”.
V-F Black-box Attack
Here we preset the performance of our black-box attack. To obtain the estimated gradient, we sample 100 random perturbation with $\epsilon=0.01$ for each data. All other settings remain identical to those used in the white-box attack. The results are shown in Figure 12 and Figure 13.
Overall, the black-box attack-implemented via zeroth-order gradient estimation closely matches the white-box performance on four datasets in terms of both metrics. As the number of known features increases, the attack performance consistently improves. When 90% of the features are known, the black-box attack achieves slightly lower performance than the white-box attack on CIFAR-10 and STL-10. This gap is more noticeable when considering TPR@0.1FPR. On Fashion-MNIST, although a more noticeable performance drop is observed, the attack remains within a practically usable range. In contrast, the difference between black-box and white-box settings on the Epsilon dataset is minimal. Additionally, it can be observed that, in the second stage of the black-box attack, using our proposed method yields a higher TPR at low FPR.
Notably, the performance degradation in the leave-one-out setting is greatly mitigated under the black-box formulation. A likely explanation is that the reconstruction problem reduces to a very low-dimensional search (one unknown feature), where finite-difference estimates act as a smoothing operator and produce more stable update directions than the raw white-box gradient. These findings indicate that zeroth-order estimation provides a practical and, in some cases, advantageous alternative to white-box gradients.
In summary, although the black-box setting results in a modest reduction in attack performance, the attack remains effective in extracting privacy information from the target model. Moreover, when considered together with partial feature availability, the black-box scenario further broadens the practical scope of membership inference attack, making it more realistic and applicable in real-world settings.
(a) CIFAR-10
(b) STL-10
(c) Fashion-MNIST
(d) Epsilon
Figure 12: Black-box attack performance (AUC) under varying known feature proportions. The x-axis represents the percentage of known features, and the y-axis shows the corresponding attack AUC. Each curve corresponds to one of the four integrated anomaly detection methods within our attack framework. We denote the case where only one feature is unknown as “loo”.
(a) CIFAR-10
(b) STL-10
(c) Fashion-MNIST
(d) Epsilon
Figure 13: Black-box attack performance (TPR@0.1FPR) under varying known feature proportions. The x-axis represents the percentage of known features, and the y-axis shows the corresponding attack TPR when TPR=0.1. Each curve corresponds to one of the four integrated anomaly detection methods within our attack framework. We denote the case where only one feature is unknown as “loo”.
@@ V-G Case Study
To simulate this scenario, we use the UCI-Diabetes dataset. We select 10 features for model training and designate race, age, and gender, common demographic attributes easily accessible to an attacker, as known features. Other experimental settings remain consistent with previous sections. The attack results are shown in Table I, demonstrating that our method can effectively compromise individual privacy in practical settings.
TABLE I: AUC performance of case study.
@@ VI Discussion
VI-A Relationship to Other Attacks
Relationship to Property Existence Attack: Property existence attack[chaudhari2023snap], as a special case of distribution inference attack[chaudhari2023snap, suri2023dissecting], aims to determine whether a given feature combination exists in the training set. The only difference from our attack lies in whether the attacker has access to the unknown features. Clearly, an attacker capable of performing PFMI can easily carry out a property existence attack, but it is difficult for a property existence attacker to perform PFMI because of the absence of some features.
Relationship to Model Inversion Attack: A common premise of existing model inversion attacks is that the adversary knows the target to be present in the training set[dibbo2023model, kahla2022label, an2022mirror, annamalai2024linear, mehnaz2022your, kabir2025disparate]. Even though the membership status of the target is unknown in the attack scenario considered in this paper, the first stage of our attack can still be regarded as a form of model inversion attack. Leveraging existing model inversion techniques thus represents a feasible approach for improving performance.
VI-B Limitation and Future Work
While our proposed method demonstrates strong attack performance, there remain several directions worth exploring in future work.
First, the effectiveness of multi-round optimization could be further improved by adopting more advanced and efficient optimization strategies, such as incorporating momentum or adaptive step-size methods. These techniques may help stabilize the reconstruction process and enhance attack performance under limited query budgets.
Second, it is important to re-evaluate the effectiveness of existing defense mechanisms in the context studied in this work. For example, machine unlearning[bourtoule2021machine, liu2025prototype] has been proposed as an effective defense against inference attacks by updating model parameters to forget a specific sample. However, even after forgetting, the model may still retain partial feature-level memorization, which could be exploited by attackers through PFMI to infer sensitive information. Other methods that have also been proven effective in defending against membership inference attacks, such as overfitting mitigation[srivastava2014dropout] and differential privacy[dwork2006differential, abadi2016deep], also require further validation in the context of this work.
@@ VII Conclusion
In this paper, we study membership inference attack under the setting where some feature values are missing. To address this scenario, we introduce a practical two-stage attack framework named MRAD which is applicable to both white-box and black-box settings. Our framework is compatible with arbitrary anomaly detection methods, enabling attackers to flexibly conduct attacks under various knowledge settings.
We conduct extensive experiments on multiple datasets to evaluate the effectiveness of our method. The results show that our framework consistently delivers strong performance across a wide range of known feature percentages. Remarkably, even when 60% of the features are missing, the attack still achieves an AUC of 0.75 on STL-10. The attack tests in the black-box scenario also demonstrate that our attack can successfully extract training data privacy, even without access to the model parameters. In addition, we use SHAP to measure feature importance and experimentally confirm that the importance of the known features has a significant impact on the attack’s performance. Finally, through a case study, we simulate a real‑world application scenario to further highlight the practical relevance of our attack method.
Our findings demonstrate that membership inference attacks can be mounted using publicly available partial data, even when private features of a sample are protected. This highlights a new dimension in privacy risk assessment.